Square root(x+2)+4=x
lease show how to solve for x

Answer by checkley79(3341)

Square root(x+2)+4=x
SQRT(X+2)=X-4 SQUARE BOTH SIDES.
X+2=X^2-8X+16
X^2-8X-X+16-2=0
X^2-9X+14=0
(X-7)(X-2)=0
X-7=0
X=7 ANS.
X-2=0
X=2 ANS.